OpenStudy (anonymous):

divide and reduce 7/12 divided by 12/16

OpenStudy (anonymous):

To divide fractions you multiply the first by the reciprocal of the second fraction: \[\frac{ 7 }{ 12 }\times \frac{ 16 }{ 12 }\] My recommendation is that you reduct the second fraction before multiplying the two fractions together to make the final simplification easier.

OpenStudy (anonymous):

so basically multiply 7/12 * 16/12. and then take the answer and reduce. I hate math.

OpenStudy (anonymous):

Simplify before multiplication it makes it easier: \[\frac{ 7 }{ 12 }\times \frac{ 16 }{ 12 }=\frac{ 7 }{ 12 }\times \frac{ 4 }{ 3 }=\frac{ 7 }{ 3 }\times \frac{ 1 }{ 3 }=\frac{ 7 }{ 9 }\]